OpenStudy (anonymous):

I really dont understand how to do this: A car dealer offers you two deals on a car that costs $16,000. Please calculate the monthly payment, given these two payment options the car dealer is offering. Payment Option 1: You can finance the car for 60 months with no interest if you make a $3,000 down payment. Payment Option 2: You can finance the car for 72 months (6 years) with 1% simple annual interest and no down payment. (Hint: To calculate simple annual interest, use the formula Interest = Principal * Rate * Time (in years). Add the amount of interest to the price of the car.) Which monthly payment amount is lower?

OpenStudy (anonymous):

Payment Option 1: $16,000 - $1,000 = $15,000 after down payment $15,000 / 60months = $250/month Payment Option 2: $16,000 * 1.02 / 72months = $226.67/month so 2nd one is better one.

OpenStudy (anonymous):

Just solve for the monthly payment options. Payment Option 1) You can finance the car for 60 months with no interest if you make a $3,000 down payment. SO: 16,000 - 3,000 = 13,000 13,000 /60 = 216.666 Monthly payment: $216.67 Payment Option 2) You can finance the car for 72 months (6 years) with 1% simple annual interest and no down payment. SO: 16,000 divided into 6 years = 2666.67 Multiply by 1% interest... 0.1 * 2666.67 = 26.67 2666.67 + 26.67 = 2693.34 per year. Divide that by 12 months. 2693.34 /12 = 224.45 Monthly Payment: $224.45

OpenStudy (anonymous):

Compare the monthly payments. Which is cheaper?

OpenStudy (anonymous):

@bloodrain its $3,000 down payment not $1,000

OpenStudy (anonymous):

dam, im sorry about tht error :/

OpenStudy (anonymous):

Monthly payment 1. Thank you so much, Jazy(:

OpenStudy (anonymous):

Payment Option 1: $16,000 - $3,000 = $13,000 after down payment $15,000 / 60months = about $216/month Payment Option 2: $16,000 * 1.02 / 72months = $226.67/month so 2nd one is better one.

OpenStudy (anonymous):

Welcome. (:

OpenStudy (anonymous):

By the way, the 0.1 was meant to be a 0.01.

OpenStudy (anonymous):

But the answer doesn't change, I just made a mistake writing it. (: